OpenStudy (anonymous):

Give the next two numbers in this pattern: 9, 4, -1, -6,..

OpenStudy (lgbasallote):

this seems like an arithmetic progression. can you find the common difference?

OpenStudy (anonymous):

A.-11, -1 B.-4, -9 C.-12, -18 those are the possible answers to the problem.

OpenStudy (experimentx):

-11, -16 ??

OpenStudy (anonymous):

yeah thats what i thought it would be! thank you!

OpenStudy (anonymous):

\[\Large d=a_n-a_{n-1} \quad , \quad a_n=a_1+(n-1)\cdot 2\] if a_1=9 then...